OpenStudy (caozeyuan):

I cant understnad why I am wrong in some questions of this problem set, question numbers are 3a and b, 8, 11, 12, 22

OpenStudy (caozeyuan):

Here is the problem set

OpenStudy (caozeyuan):

sorry, should be11, 13, 22

OpenStudy (mathmale):

Please choose one problem and get started on it. You might want to list what you know already and what ou believe you'll need to progress further. I'd be glad to give you feedback on your own efforts. By the way, I cannot read your attachment except by saving it to Word and then opening it in Word.

OpenStudy (caozeyuan):

3a and 3b I have wrote 1/2*1/2*1/6=1/24, but wrong, why? and I got c correct?!

OpenStudy (caozeyuan):

In 11, I use binomial distrubution e.g. for a I have 10C0*(4/5)^10=.107

OpenStudy (caozeyuan):

in 8 I did 36C4, so should I do 36*35*34*33? why?

OpenStudy (caozeyuan):

nvm, I got 13!

OpenStudy (caozeyuan):

anybody? any idea?

OpenStudy (caozeyuan):

please, exam on thursday, need answer now!

OpenStudy (anonymous):

sorry i cant help but to make u feel better heres a medal

OpenStudy (caozeyuan):

thx

OpenStudy (caozeyuan):

@Photon336

OpenStudy (photon336):

@zepdrix